OpenStudy (anonymous):

[ d/dx]( ( 4√{x} ) ln(x)) = ? 4 is the upper on the square root so it means 1/4

OpenStudy (xishem):

Alright. So you'll just need to use the product rule, right?

OpenStudy (anonymous):

But I have no idea how to do it..

OpenStudy (xishem):

OK, do you know what the product rule is?

OpenStudy (anonymous):

Yea. Oh wait the question is [ d/dx]( ( 4√{x} )^(ln(x))) so it should be power rule

OpenStudy (xishem):

You actually have to use both the power rule and the product rule:\[\frac{d}{dx}(\sqrt[4]{x}\ln(x))=\frac{d}{dx}(\sqrt[4]{x})*\ln(x)+\sqrt[4]{x}*\frac{d}{dx}(\ln(x))\]Does that help?

OpenStudy (anonymous):

So is it 1/4x^(-3/4)(ln(x))+x^(3/4).?

OpenStudy (xishem):

The right term should be this, I believe:\[x^{1/4}*\frac{1}{x}=x^{-3/4}\]

OpenStudy (anonymous):

1/4x^(-3/4)(ln(x))+x^(-3/4.? But it is still incorrect...

OpenStudy (xishem):

No, that's right: http://www.wolframalpha.com/input/?i=d%2Fdx%28x%5E%281%2F4%29*ln%28x%29%29

OpenStudy (anonymous):

Um...? The web work said wrong...

OpenStudy (anonymous):

Wait... The question is to the power of ln(x)

OpenStudy (xishem):

Write it out?